# Class-Based Expurgation Attains Csiszár’s Expurgated Source-Channel Exponent

### v1

@@ I Introduction
We study the transmission of non-equiprobable messages from a discrete memoryless source (DMS) with distribution $P^{k}(\boldsymbol{v})=\prod_{i=1}^{k}P_{V}(v_{i})$, where $\boldsymbol{v}=(v_{1},...,v_{k})\in\mathcal{V}^{k}$ is the source message, and $\mathcal{V}$ is a finite discrete alphabet. The channel is a discrete memoryless channel (DMC) given by $W^{n}(\boldsymbol{y}|\boldsymbol{x})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, $\boldsymbol{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$ and $\boldsymbol{y}=(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$, where $\mathcal{X}$ and $\mathcal{Y}$ are discrete alphabets with cardinalities $|\mathcal{X}|$ and $|\mathcal{Y}|$, respectively. An encoder maps the length-$k$ source message $\boldsymbol{v}$ to a length-$n$ codeword $\boldsymbol{x}_{\boldsymbol{v}}$, which is then transmitted over the channel. We refer to $t\triangleq k/n$ as the transmission rate. Based on the channel output $\boldsymbol{y}$ the decoder guesses which source message was transmitted.
@@
Expurgation has received less attention than random coding bounds in JSCC. In [4], Csiszár derived two expurgated exponents, which were later recovered by Scarlett [5] using Gallager’s classical expurgation method [6, Sec. 5.7]. Motivated by the preceding discussion, we study expurgation under a partition of the source messages. We show that, as in the random-coding case, two carefully chosen classes are sufficient to achieve Csiszár’s second expurgated exponent [4].
@@
The number of source type classes in the set $\mathcal{V}^{k}$ is denoted by $N_{k}$, which grows polynomially with $k$ [7, Lemma 2.2]. Throughout, we use the indices $i$ and $j$ to refer to source types, and $c$ to indicate the class to which a source message belongs under a given partitioning. The total number of classes in the partitioning is denoted by $m$. Moreover, for a source type $P_{i}$, we define $R_{i}\triangleq tH(P_{i})$.
@@ II Csiszár’s Expurgated Error Exponents
In this section, we introduce definitions and summarize existing expurgated results in JSCC that are relevant to this work. The source reliability function [7, Ch. 9] is given by
@@
Using the packing lemma [1, Lemma 6], Csiszár [4] derived two achievable expurgated exponents for joint source-channel coding. The first exponent was derived by assigning the same composition to all source messages, in a similar way to Gallager’s [6, Problem 5.16] for the i.i.d. ensemble. The expression is given by
@@
It is not known which of the two exponents is higher. Csiszár partitions the source sequences according to their type (i.e., $m=N_{k}$), assigning a distinct codeword distribution to each one. A more general partitioning was later proposed in [3], where they derived the maximum a posteriori (MAP) random-coding error exponent for arbitrary partitioning of the source sequence set. It was further shown that a specific two-class partitioning suffices to recover Csiszár’s random-coding exponent. In this work, we consider a less general partitioning, where each class contains one or more full source type classes. This setting includes Csiszár’s construction as a special case. We then show that two carefully chosen classes are sufficient to achieve $E_{J,2}^{\mathrm{ex}}(t,P_{V})$.
The work in [5] revisits expurgation in joint source-channel coding and rederives Csiszár’s exponents from [4] using Gallager’s classical expurgation method [6, Sec. 5.7]. The key result from[5] is summarized in the following lemma.
@@
where $\rho_{ij}>0$, and $p_{e}\big{(}\boldsymbol{v},j,\mathcal{C}\big{)}$ refers to the probability that, given $\boldsymbol{V}=\boldsymbol{v}$, there exists some $\boldsymbol{\bar{v}}\in\mathcal{T}^{k}(P_{j})$ that yields a decoding metric at least as high as that of $\boldsymbol{v}$. Moreover, the expectation is taken with respect to the given (not necessarily distinct) codeword distributions $\{P_{\boldsymbol{X}}^{(j)}\}_{j=1}^{N_{k}}$.
@@ III-A Primal-Domain
Let $\mathcal{A}_{1},...,\mathcal{A}_{m}$ be a given arbitrary partition of the source sequence, each associated with a codeword distribution $Q_{c}\in\mathcal{P}_{n}(\mathcal{X})$. Then there exists a codebook such that
@@
In the next theorem, we consider the partitioning that maximizes the exponent in Lemma 2.
@@ III-B Dual-Domain
The next result extends [2, Lemma 3] to the expurgated exponent for a given set of codeword distributions.
@@
The function $\overline{E}_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$ is the concave hull on the interval $[1,\infty)$ of the function $E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$.
@@
If the partitioning is based on source type classes and the associated distributions $\{Q_{i}\}_{i=1}^{N_{k}}$ are chosen according to
@@ III-C Two-Class Partition of the Source Sequences
The following corollary is the expurgated analogue of [3, Corollary 1]. As in the random-coding case, a direct application of Carathéodory’s theorem [11, Cor. 17.1.5] implies that any point on the graph of $\overline{E}_{\mathrm{x}}^{\prime}(\lambda)$ can be expressed as a convex combination of two points on the graph of ${E_{\mathrm{x}}^{\prime}}(\lambda)$. As a result, there exists a pair of distributions $\{Q,Q^{\prime}\}$ such that both points lie on the graph of ${E_{\mathrm{x}}^{\prime}}(\lambda,\{Q,Q^{\prime}\})$. Optimizing over all such pairs gives the following result.
@@
There exists a partition of the source sequences into two classes, each associated with a distribution from a given pair, such that $E_{J,2}^{\mathrm{ex}}(t,P_{V})$ in (10) is achievable.

### v2

@@ I Introduction
[…]
@@
Expurgation has received less attention than random coding bounds in JSCC. In [4], Csiszár derived two expurgated exponents, which were later recovered by Scarlett [5] using a new method based on source-type duplication and type-by-type expurgation.
Motivated by the preceding discussion, we study expurgation under a partition of the source messages. We show that, as in the random-coding case, two carefully chosen classes are sufficient to achieve Csiszár’s second expurgated exponent [4].
@@
The number of source type classes in the set $\mathcal{V}^{k}$ is denoted by $N_{k}$, which grows polynomially with $k$ [6, Lemma 2.2]. Throughout, we use the indices $i$ and $j$ to refer to source types, and $c$ to indicate the class to which a source message belongs under a given partitioning. The total number of classes in the partitioning is denoted by $m$. Moreover, for a source type $P_{i}$, we define $R_{i}\triangleq tH(P_{i})$.
@@ II Csiszár’s Expurgated Error Exponents
In this section, we introduce definitions and summarize existing expurgated results in JSCC that are relevant to this work. The source reliability function [6, Ch. 9] is given by
@@
Using the packing lemma [1, Lemma 6], Csiszár [4] derived two achievable expurgated exponents for joint source-channel coding. The first exponent was derived by assigning the same composition to all source messages, in a similar way to Gallager’s [9, Problem 5.16] for the i.i.d. ensemble. The expression is given by
@@
It is not known which of the two exponents is higher. Csiszár partitions the source sequences according to their type (i.e., $m=N_{k}$), assigning a distinct codeword distribution to each one. A more general partitioning was later proposed in [3], where they derived the maximum a posteriori (MAP) random-coding error exponent for arbitrary partitioning of the source sequence set. It was further shown that a specific two-class partitioning suffices to recover Csiszár’s random-coding exponent. The work in [10] considers a less general partitioning, where each class contains one or more full source type classes, and shows that this recovers the MAP random-coding exponent of [3] using the universal maximum mutual information (MMI) decoder. In this work, we adopt the partitioning of [10], and show that two carefully chosen classes are also sufficient to achieve $E_{J,2}^{\mathrm{ex}}(t,P_{V})$.
The work in [5] revisits expurgation in joint source-channel coding and rederives Csiszár’s exponents from [4] using a new expurgation method. The key result from[5] is summarized in the following lemma.
@@
where $\rho_{ij}>0$, and $p_{e}\big{(}\boldsymbol{v},j,\mathsf{C}\big{)}$ refers to the probability that, given $\boldsymbol{V}=\boldsymbol{v}$, there exists some $\boldsymbol{\bar{v}}\in\mathcal{T}^{k}(P_{j})$ that yields a decoding metric at least as high as that of $\boldsymbol{v}$. Moreover, the expectation is taken with respect to the given (not necessarily distinct) codeword distributions $\{P_{\boldsymbol{X}}^{(j)}\}_{j=1}^{N_{k}}$.
@@ III-A Primal-Domain
Let $\mathcal{A}_{1},...,\mathcal{A}_{m}$ be a given arbitrary partition of the source sequences, each associated with a codeword distribution $Q_{c}\in\mathcal{P}_{n}(\mathcal{X})$. Then there exists a codebook such that
@@
The next theorem considers the partitioning that maximizes the exponent in Lemma 2, for the given codeword distributions.
@@ III-B Dual-Domain
Theorem 2 is the dual counterpart of Theorem 1, and their equivalence follows from Lagrange duality. The following result extends [2, Lemma 3] to the expurgated exponent for a given set of codeword distributions.
@@
The function $\overline{E}_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$ is the concave hull on the interval [1,)[1$\infty$)[ 1 , ∞ ) of the function $E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$.
@@
If the partitioning is based on source type and the associated distributions $\{Q_{i}\}_{i=1}^{N_{k}}$ are chosen according to
@@ III-C Two-Class Partition of the Source Sequences
The following corollary is the expurgated analogue of [3, Corollary 1]. As in the random-coding case, a direct application of Carathéodory’s theorem [12, Cor. 17.1.5] implies that for any $\lambda_{0}$, the value $\overline{E}_{\mathrm{x}}^{\prime}(\lambda_{0})$ can be exprssed as a convex combination of two points on the graph of $E_{\mathrm{x}}^{\prime}(\lambda)$. Let these points correspond to $\lambda_{1}$ and $\lambda_{2}$, and let $Q_{1}$ and $Q_{2}$ be the respective maximisers of $E_{\mathrm{x}}^{\prime}(\tilde{Q},\lambda)$ at $\lambda_{1}$ and $\lambda_{2}$. It follows that both points also lie on the graph of $E_{\mathrm{x}}^{\prime}(\{Q_{1},Q_{2}\},\lambda)$, and thus $\overline{E}_{\mathrm{x}}^{\prime}(\lambda_{0})$ also lies on its concave hull. This shows that a two-class partitioning suffices to achieve $\overline{E}_{\mathrm{x}}^{\prime}(\lambda_{0})$. Optimizing over all such pairs gives the following result.
@@
There exists a partition of the source sequences into two classes, such that $E_{J,2}^{\mathrm{ex}}(t,P_{V})$ in (10) is achievable.